OpenStudy (anonymous):

find the linear functions. the x and y intercepts of the inverse function are 5 and -1 respectively

OpenStudy (saifoo.khan):

a bit tricky, but easy! :D

OpenStudy (anonymous):

What does having the respectively do? or is this one of those words just to throw you off.

OpenStudy (saifoo.khan):

you have to find the equation of the inverse first, then find it's inverse, finally that will be your answer.

OpenStudy (anonymous):

wow

OpenStudy (saifoo.khan):

yo? stuck?

OpenStudy (anonymous):

yea

OpenStudy (saifoo.khan):

u just need the answers right?

OpenStudy (anonymous):

no I have to show my work

OpenStudy (saifoo.khan):

\[y = \frac x5-1\] now find it's inverse.

OpenStudy (saifoo.khan):

Reallly? :/

OpenStudy (anonymous):

\[y=x/5-1=4\]

OpenStudy (saifoo.khan):

HOW DID YOU GOT THAT?

OpenStudy (anonymous):

i don't know

OpenStudy (saifoo.khan):

what should i do now? :o

OpenStudy (anonymous):

pull your hair out i guess

OpenStudy (anonymous):

I know be patience(?) with me

OpenStudy (saifoo.khan):

Lol, Dont worry man!! im patient. im just asking.. what u want now? working step by step?

OpenStudy (anonymous):

yes that would be great

OpenStudy (saifoo.khan):

(5,0) and (0,-1) slope = (-1,-0)/(0-5) slope = 1/5 y = mx + b 0 = -1(5)/5 + b b = -1 So equation for inverse = y = 1x/5 -1 now finding it's inverse, y = (x-5)/5 5y + 5 = x \[f^{-1}(x)=5x+5\]

OpenStudy (saifoo.khan):

So the equation for the final answer is, y = 5x + 5

OpenStudy (anonymous):

ok I understand how to find the inverse now I guess i just needed to see it done.

OpenStudy (anonymous):

so any time that i only have 2 numbers I will stick a 0 in there to help out.

OpenStudy (saifoo.khan):

No!!! remember, when x intercepts y = 0 when y intercepts x = 0

OpenStudy (anonymous):

thats right Thanks

OpenStudy (saifoo.khan):

You welcome.